OpenStudy (anonymous):

how do you find the sum of the series sigma n=0 ; 4 (cosn*pi) (3/4)^n

OpenStudy (anonymous):

use the equation maker

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} 4 (\cos n \pi) (\frac{ 3 }{ 4 })^{n}\]

OpenStudy (anonymous):

are you sure it is till infinity?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

Perhaps you can use some trig identities in conjunction with the ratio test.